OpenStudy (anonymous):

8 2/17* 1 15/17 = a.16 b.24 c.8 d.12

OpenStudy (dtan5457):

\[8\frac{ 2 }{ 17 }\times 1\frac{ 15 }{ 17 }\]

OpenStudy (dtan5457):

like this?

OpenStudy (dtan5457):

@lissy44

OpenStudy (anonymous):

yes

OpenStudy (dtan5457):

Can't be it. None of your answer choices work.

OpenStudy (dtan5457):

Please check your question again

OpenStudy (anonymous):

2/17 * 15/17 2*15=30 30/17, put it into proper fraction form and multiply 8*1